OpenStudy (anonymous):

Solve the equation by completing the square X2-12X=-28 Please show all of your work

OpenStudy (anonymous):

Need to set equal to zero first

OpenStudy (anonymous):

\[x^2-12x+28=0\]

OpenStudy (anonymous):

okk...

OpenStudy (anonymous):

now we take half of the center term and write this (x - 6)^2. Half of the center term is -6. Next we have to take out what we put in. We put in -6^2=36 so we need to take this out. So we add (x-6)^2 - 36 + 28. After simplifying we get...\[(x-6)^2-8\]

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

Yeahh i think i get it..

OpenStudy (anonymous):

but then what do u do next

OpenStudy (anonymous):

like u go ahead and take the square root of that?

OpenStudy (anonymous):

O ya I guess we still have to solve it. What I showed was only completing the square.\[(x-6)^2-8=0\]\[(x-6)^2=8\]There will be 2 answers because (x-6) can be positive or negative since when it is squared it will become positive. I do think you should take the square root of each side. then solve for x for the positive square root of 8 and the negative square root of 8

OpenStudy (anonymous):

mmmm ok thankss